# Combinatorial flats and Schubert varieties of subspace arrangements

### v2

@@ 1. Introduction
To remedy these pathologies, we introduce the lattice of combinatorial flats $\mathcal{L}_{P}$ of a polymatroid. When $P$ is a matroid, $\mathcal{L}_{P}$ is the usual lattice of flats. When $P$ is realized by a subspace arrangement $V_{1},\ldots,V_{N}$, the lattice $\mathcal{L}_{P}$ is the poset of flats of a collection of general flags with minimal elements $V_{1},\ldots,V_{N}$ (see 1.2).
@@
We define $\mathcal{L}_{P}$ to be the set of combinatorial flats, ordered by inclusion. We will see in Section 3.4 that $\mathcal{L}_{P}$ does not depend on $\mathbf{n}$ (although the underlying set of combinatorial flats does; we write $\mathcal{L}_{P,\mathbf{n}}$ when we wish to emphasize this).
@@
Suppose that $V_{1},V_{2},V_{3},V_{4}\subseteq V=\mathbb{C}^{3}$ is a linear subspace arrangement, where $V_{1},V_{2},$ and $V_{3}$ are two dimensional, $V_{1}\cap V_{2}\cap V_{3}$ is one dimensional, and $V_{4}$ is one dimensional and not contained in $V_{i}$ for $i=1,2,3$. This arrangement defines the polymatroid $P$ on $E=\{1,2,3,4\}$ with $\operatorname{rk}_{P}(S)=3-\dim\cap_{i\in S}V_{i}$. Flats of $P$ correspond to intersections of $V_{i}$’s. The lattice of flats is neither graded nor top-heavy (Fig. 1).
@@
12341231234$\emptyset$
Figure 1. Left: the projectivization of the subspaces arrangement $V_{1},\ldots,V_{4}$ from 1.2 (solid), along with the projectivization of a general plane $V_{4,1}$ containing $V_{4}$ (dashed). Right: the lattice of flats of the polymatroid $P$ realized by $\{V_{1},V_{2},V_{3},V_{4}\}$.
On the other hand, $\mathcal{L}_{P}\cong\mathcal{L}_{P,\mathbf{n}}$ (with $\mathbf{n}=(1,1,1,2)$) is both graded and top-heavy (Fig. 2). Elements of $\mathcal{L}_{P}$ correspond to flats of the arrangement $\{V_{1},V_{2},V_{3},V_{4},V_{4,1}\}$, where $V_{4,1}$ is a general plane containing $V_{4}$. There is a lattice homomorphism embedding the lattice of flats of $P$ into $\mathcal{L}_{P}$. This construction is explained in full generality in 3.7.
@@
Figure 2. The lattice of combinatorial flats of $P$ from 1.2. Join-irreducibles are underlined. Blue elements are in the image of the embedding of $P$’s lattice of flats into $\mathcal{L}_{P}$.
@@
Lattices of combinatorial flats can be axiomatized in a manner that closely resembles the well-known definition of geometric lattices (3.20). Consequently, they constitute a new cryptomorphic definition of simple polymatroids.
@@
Work of Hassett-Tschinkel [HT99] (reviewed in Section 4.1) shows that there exists action of the additive group $\mathbb{K}^{\mathbf{n}}:=\mathbb{K}^{n_{1}}\times\cdots\times\mathbb{K}^{n_{N}}$ on $\mathbb{P}^{\mathbf{n}}:=\mathbb{P}^{n_{1}}\times\cdots\times\mathbb{P}^{n_{N}}$, such that $\mathbb{P}^{\mathbf{n}}$ is partitioned into finitely many orbits $O_{\mathbf{s}}$, corresponding to multisets $\mathbf{s}\leq\mathbf{n}$. Under the assumption that $V$ is “polymatroid general”, a mild genericity condition described in Section 4.6 that constrains the intersection of $V$ with certain coordinate subspaces, we define $Y_{V}$ to be the $V$-orbit closure of the origin in $\mathbb{P}^{\mathbf{n}}$ under the Hasset-Tschinkel action of $\mathbb{K}^{\mathbf{n}}$ on $\mathbb{P}^{\mathbf{n}}.$
@@
    $[\overline{Y_{V}\cap O_{\mathbf{s}}}]=\sum_{\mathbf{b}\text{ basis of $\mathbf{s}$}}c_{\mathbf{b}}[\mathbb{P}^{\mathbf{b}}],$
@@ 1.4. Organization
We review polymatroids, multisymmetric matroids, and operations on them in Section 2, before commencing study of combinatorial flats in Section 3.
We construct combinatorial flats in Section 3.1, then consider the effects of deletion and truncation in Section 3.2. These sections are the only portions of Section 3 required for Section 4. In the remainder of Section 3, we work up to a proof of 1.1 (Sections 3.3 and 3.4) and provide axioms for $\mathcal{L}_{P}$ (Section 3.5)
Section 4 concerns Schubert varieties of subspace arrangements. Foundational work of Hassett and Tschinkel [HT99] is reviewed in Sections 4.1 and 4.2. We construct $Y_{V}$ in Section 4.4, then compute the support of its Chow class in Section 4.5. In Section 4.6, we discuss “polymatroid generality”, a condition on $V$ that guarantees $Y_{V}$ consists of finitely many $V$-orbits. Assuming this condition, we identify strata of $Y_{V}$ in Section 4.7, culminating with a proof of 1.3. In Section 4.8, we prove 1.7.
At the last, we pose a large number of questions about Schubert varieties of subspace arrangements in Section 5. Both algebraic and geometric aspects of Schubert varieties of subspace arrangements seem markedly more complicated than those of Schubert varieties of hyperplane arrangements, yet examples suggest a good theory lies in wait. Interested parties are encouraged to seek their fortune.
@@ 1.6. Acknowledgements
We thank Lucas Gierczak for helping to sort out the relationship between the “matricubes” of [AG24] and the objects of the present work (see 3.8), and Matt Larson for helpful comments on our draft. We are all partially supported by the Simons Foundation. The first author is partially supported by NSF grant DMS-2039316, and the last author is partially supported by NSF grant DMS-1926686. We are grateful to the Institute for Advanced Study (IAS). The second and third authors also thank the Beijing International Center for Mathematical Research (BICMR) for their hospitality and support during the preparation of this paper.
@@ 2. Polymatroids
{caged polymatroids ($\{\text{caged polymatroids $(P,\mathbf{n})$}\}$$\overset{\sim}{\longrightarrow}\{\text{multisymmetric matroids on$\widetilde{E}$}\}$
@@
The matroid lift satisfies $\operatorname{rk}_{P}(S)=\operatorname{rk}_{\widetilde{P}}(\cup_{i\in S}E_{i})$. By applying this fact and 2.4, we learn that the closure in $\widetilde{P}$ of $\cup_{i\in S}E_{i}$ is $\cup_{i\in\overline{S}}E_{i}$, where $\overline{S}$ is the closure of $S$ in $P$.
@@
By 2.5, $\widetilde{P}$ is the unique multisymmetric matroid on $\widetilde{E}$ that lifts $P$.
@@
$\operatorname{rk}_{\widetilde{P}}(A)=\operatorname{rk}_{\widetilde{P}}(\overline{A})$ $=|\overline{A}\setminus{\overline{A}}^{\mathrm{geo}}|+\operatorname{rk}_{\widetilde{P}}(\overline{A}^{\mathrm{geo}})$ by 2.4
@@
[…]
@@
With notation as in 2.2, if $P$ is realized by $V$, then $P\setminus A$ is realized by $\pi_{E\setminus A}(V)$.
@@
$\operatorname{rk}_{T_{\cup_{i\in S}E_{i}}\widetilde{P}}(\cup_{i\in A}E_{i})$ =$=\begin{cases}\operatorname{rk}_{\widetilde{P}}(\cup_{i\in A}E_{i})-1,&\text{if $\operatorname{rk}_{\widetilde{P}}(\cup_{i\in A}E_{i})=\operatorname{rk}_{\widetilde{P}}(\cup_{i\in A\cup S}E_{i})$}\\ \operatorname{rk}_{\widetilde{P}}(\cup_{i\in A}E_{i}),&\text{otherwise}\end{cases}$
=$=\begin{cases}\operatorname{rk}_{P}(A)-1,&\text{if $\operatorname{rk}_{P}(A)=\operatorname{rk}_{P}(A\cup S)$}\\ \operatorname{rk}_{P}(A),&\text{otherwise}\end{cases}$
@@
Consequently, $\widetilde{T_{S}P}=T_{\cup_{i\in S}E_{i}}\widetilde{P}$ by 2.5. ∎
@@
With notation as in 2.2, if $S\subset E$, then $T_{S}P$ is realized by $V\cap\pi^{-1}_{E\setminus S}(H)$, with $H$ a general hyperplane in $\prod_{i\in S}W_{i}$.
@@
We may think of both matroids as being multisymmetric on $\widetilde{E}\setminus j$. Geometric subsets of $\widetilde{E}\setminus j$ are of the form $(\cup_{k\in A}E_{k})\setminus j$, with $A\subset E$. By 2.5 and 2.7, it suffices to show that
@@
Otherwise, suppose that $\operatorname{rk}_{P}(A)<\operatorname{rk}_{P}(A\setminus i)+\operatorname{rk}_{P}(i)$. In this case, 2.4 implies that $\overline{(\cup_{k\in A}E_{k})\setminus j}\supset\cup_{k\in A}E_{k}$ in $\widetilde{P}$. Hence,
@@
We defer discussion of realizing reductions to 3.17, where we will see it in action en route to proving 1.1(ii).
@@ 3. Combinatorial flats
We introduce the poset of combinatorial flats of a polymatroid $P$. We construct the poset using a chosen lift $\widetilde{P}$ in Section 3.1, then record the effects of deletion and truncation in Section 3.2. In Sections 3.3 and 3.4, we establish order-theoretic properties and consider the effects of reduction, culminating in a proof of 1.1. Finally, in Section 3.5, we axiomatize posets of combinatorial flats. This provides a cryptomorphic definition of simple polymatroids that generalizes the cryptomorphism of simple matroids and geometric lattices.
Our results on Schubert varieties of subspace arrangements depend only on Sections 3.1 and 3.2, so geometrically-inclined readers may go directly to Section 4 after reading these.
Throughout Section 3, we fix notations:
@@
  • $\operatorname{rk}_{P}(\mathbf{s})$ is the rank of a multiset $\mathbf{s}$, in the sense of Section 1.2.
@@
A combinatorial flat of $(P,\mathbf{n})$ is a multiset $\mathbf{s}\in\mathbb{Z}_{\geq 0}^{E}$ such that there exists a flat $F$ of $\widetilde{P}$ satisfying $\mathbf{s}=(|F\cap E_{1}|,\ldots,|F\cap E_{N}|)$. (Rephrased in the language of Section 1.5, $\mathbf{s}$ is a combinatorial flat if it is represented by a flat of $\widetilde{P}$.)
@@
Viewing $\mathcal{L}_{P,\mathbf{n}}$ as a quotient and applying 2.4, we learn that rank and closure of multisets have the following properties:
@@
Choose $S\subset\widetilde{E}$ representing $\mathbf{s}$. If $\mathbf{b}$ is a basis of $\mathbf{s}$, then $\mathbf{b}$ is represented by a basis $B$ of $S$ in $\widetilde{P}$ by 2.9. The first assertion holds because $\operatorname{rk}_{P,\mathbf{n}}(\mathbf{s})=\operatorname{rk}_{\widetilde{P}}(S)=|B|=|\mathbf{b}|$.
If $\mathbf{s}$ is a combinatorial flat, then $S\setminus S^{\mathrm{geo}}$ consists of coloops of the flat $S$ by 2.4. Hence, bases of $S$ are of the form $B^{\prime}\cup(S\setminus S^{\mathrm{geo}})$ with $B^{\prime}$ a basis for $S^{\mathrm{geo}}$. The second assertion follows. ∎
@@
3.4 shows that all bases of a multiset $\mathbf{s}$ have the same cardinality, so $\operatorname{rk}_{P}(\mathbf{s})$, as defined in Section 1.2, is well-defined.
@@
3.4 explains the equivalence of the different definitions of “rank” and “combinatorial flat” given in Section 1.2 and Section 3.1. The first statement of 3.4 says exactly that $\operatorname{rk}_{P,\mathbf{n}}(\mathbf{s})=\operatorname{rk}_{P}(\mathbf{s})$ for any multiset $\mathbf{s}$. When $\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$, this means $\operatorname{rk}_{P}(\mathbf{s})$ is also equal to the rank of $\mathbf{s}$ in the graded poset $\mathcal{L}_{P,\mathbf{n}}$. We will henceforth write only $\operatorname{rk}_{P}(\mathbf{s})$.
For equivalence of the two definitions of “combinatorial flat”, choose $S\subset\widetilde{E}$ representing $\mathbf{s}$. If $s_{i}<n_{i}$, then $\mathbf{s}+\mathbf{e}_{i}$ is represented $S\cup e$ for any $e\in E_{i}\setminus S$. In the sense of Section 1.2, $\mathbf{s}$ is a combinatorial flat if and only if
@@
for all $e\in E_{i}$ with $s_{i}<n_{i}$. This is precisely the statement that $S$ is a flat of $\widetilde{P}$, which is the notion of “combinatorial flat” defined in Section 3.1.
@@
Suppose that the polymatroid $P$ is realized by a subspace arrangement $V_{1},\ldots,~{}V_{N}~{}\subset~{}V$. Rephrasing 2.10 in the language of arrangements, if $\mathbf{n}$ is a cage for $P$, then $\widetilde{P}$ is realized by any arrangement $\{H_{ij}:1\leq i\leq N,1\leq j\leq n_{i}\}$ in which $H_{i1},\ldots,H_{in_{i}}$ are general hyperplanes containing $V_{i}$. The lattice of flats $\mathcal{L}_{\widetilde{P}}$ has an action by a product of symmetric groups $\mathfrak{S}$, and the quotient $\mathcal{L}_{\widetilde{P}}/\mathfrak{S}$ is isomorphic to $\mathcal{L}_{P}$ (see also Section 3.1).
@@
From the flag perspective of 3.7, the difference between matricubes and combinatorial flats is that matricubes model all collections of flags, while combinatorial flats model only collections of general flags. Combinatorially, this difference manifests in several ways: combinatorial flats obey 3.2, are top-heavy, and admit a reasonable notion of basis, but none of these properties hold for matricubes in general. (Failure of top-heaviness and 3.2 for matricubes can be seen in the right-hand example of [AG24, Section 3.1]. The difficulties of bases for matricubes are discussed in [AG24, Section 9.1].)
@@
The if direction follows from 2.12 and 2.13. For the “only if”, suppose that $\pi_{E\setminus i}(\mathbf{s})$ is a combinatorial flat of $(P\setminus i,\pi_{E\setminus i}(\mathbf{n}))$. We will show that $\mathbf{s}$ must be a flat by proving $\mathbf{s}=\overline{\mathbf{s}}$. By 3.2 and the hypothesis that $\overline{s}_{i}<n_{i}$, it is always the case that $s_{i}=\overline{s}_{i}$. Hence, it suffices to show that $s_{j}=\overline{s}_{j}$ for $j\neq i$.
By 2.12 and 2.13, there is a combinatorial flat $\mathbf{s}^{\prime}$ of $(P,\mathbf{n})$ with $\pi_{E\setminus i}(\mathbf{s}^{\prime})=\pi_{E\setminus i}(\mathbf{s})$. If $s^{\prime}_{i}=n_{i}$, then $\mathbf{s}<\mathbf{s}^{\prime}$, so $\mathbf{s}\leq\overline{\mathbf{s}}<\mathbf{s}^{\prime}$. Entrywise, this implies $s_{j}\leq\overline{s}_{j}\leq s_{j}^{\prime}=s_{j}$, which gives the desired equality.
Otherwise, $s_{i}^{\prime}<n_{i}$. By Lemmas 3.2 and 3.3, $\mathbf{s}^{\prime\prime}:=\mathbf{s}^{\prime}-s_{i}^{\prime}\mathbf{e}_{i}=\mathbf{s}-s_{i}\mathbf{e}_{i}$ is a combinatorial flat of rank at least $\operatorname{rk}(\mathbf{s})-s_{i}$. By 3.2, $\overline{\mathbf{s}}-s_{i}\mathbf{e}_{i}$ is also a combinatorial flat, and has rank $\operatorname{rk}(\overline{\mathbf{s}})-s_{i}$. Since $\mathbf{s}^{\prime\prime}\leq\overline{\mathbf{s}}-s_{i}\mathbf{e}_{i}$, these two combinatorial flats must be equal. In particular, for $j\neq i$, $s_{j}^{\prime\prime}=\overline{s}_{j}$. Since $s_{j}^{\prime\prime}=s_{j}$ by construction, this means $s_{j}=\overline{s}_{j}$, as desired. ∎

### v3

@@ 1. Introduction
To remedy these pathologies, we introduce the lattice of combinatorial flats $\mathcal{L}_{P}$ of a polymatroid. When $P$ is a matroid, $\mathcal{L}_{P}$ is the usual lattice of flats. When $P$ is realized by a subspace arrangement $V_{1},\ldots,V_{N}$, the lattice $\mathcal{L}_{P}$ is the poset of flats of a collection of general flags with minimal elements $V_{1},\ldots,V_{N}$ (see ˜1.2).
@@
We define $\mathcal{L}_{P}$ to be the set of combinatorial flats, ordered by inclusion. We will see in Section˜3.4 that $\mathcal{L}_{P}$ does not depend on $\mathbf{n}$ (although the underlying set of combinatorial flats does; we write $\mathcal{L}_{P,\mathbf{n}}$ when we wish to emphasize this).
@@
Suppose that $V_{1},V_{2},V_{3},V_{4}\subseteq V=\mathbb{C}^{3}$ is a linear subspace arrangement, where $V_{1},V_{2},$ and $V_{3}$ are two dimensional, $V_{1}\cap V_{2}\cap V_{3}$ is one dimensional, and $V_{4}$ is one dimensional and not contained in $V_{i}$ for $i=1,2,3$. This arrangement defines the polymatroid $P$ on $E=\{1,2,3,4\}$ with $\operatorname{rk}_{P}(S)=3-\dim\cap_{i\in S}V_{i}$. Flats of $P$ correspond to intersections of $V_{i}$’s. The lattice of flats is neither graded nor top-heavy (Fig.˜1).
$V_{2}$$V_{1}$$V_{3}$$V_{4}$$V_{4,1}$
@@
Figure 1. Left: the projectivization of the subspaces arrangement $V_{1},\ldots,V_{4}$ from ˜1.2 (solid), along with the projectivization of a general plane $V_{4,1}$ containing $V_{4}$ (dashed). Right: the lattice of flats of the polymatroid $P$ realized by $\{V_{1},V_{2},V_{3},V_{4}\}$.
On the other hand, $\mathcal{L}_{P}\cong\mathcal{L}_{P,\mathbf{n}}$ (with $\mathbf{n}=(1,1,1,2)$) is both graded and top-heavy (Fig.˜2). Elements of $\mathcal{L}_{P}$ correspond to flats of the arrangement $\{V_{1},V_{2},V_{3},V_{4},V_{4,1}\}$, where $V_{4,1}$ is a general plane containing $V_{4}$. There is a lattice homomorphism embedding the lattice of flats of $P$ into $\mathcal{L}_{P}$. This construction is explained in full generality in ˜3.7.
@@
Figure 2. The lattice of combinatorial flats of $P$ from ˜1.2. Join-irreducibles are underlined. Blue elements are in the image of the embedding of $P$’s lattice of flats into $\mathcal{L}_{P}$.
@@
Lattices of combinatorial flats can be axiomatized in a manner that closely resembles the well-known definition of geometric lattices (˜3.20). Consequently, they constitute a new cryptomorphic definition of simple polymatroids.
@@
Work of Hassett-Tschinkel [HT99] (reviewed in Section˜4.1) shows that there exists action of the additive group $\mathbb{K}^{\mathbf{n}}:=\mathbb{K}^{n_{1}}\times\cdots\times\mathbb{K}^{n_{N}}$ on $\mathbb{P}^{\mathbf{n}}:=\mathbb{P}^{n_{1}}\times\cdots\times\mathbb{P}^{n_{N}}$, such that $\mathbb{P}^{\mathbf{n}}$ is partitioned into finitely many orbits $O_{\mathbf{s}}$, corresponding to multisets $\mathbf{s}\leq\mathbf{n}$. Under the assumption that $V$ is “polymatroid general”, a mild genericity condition described in Section˜4.6 that constrains the intersection of $V$ with certain coordinate subspaces, we define $Y_{V}$ to be the $V$-orbit closure of the origin in $\mathbb{P}^{\mathbf{n}}$ under the Hasset-Tschinkel action of $\mathbb{K}^{\mathbf{n}}$ on $\mathbb{P}^{\mathbf{n}}.$
@@
[…]
@@ 1.4. Organization
We review polymatroids, multisymmetric matroids, and operations on them in Section˜2, before commencing study of combinatorial flats in Section˜3.
We construct combinatorial flats in Section˜3.1, then consider the effects of deletion and truncation in Section˜3.2. These sections are the only portions of Section˜3 required for Section˜4. In the remainder of Section˜3, we work up to a proof of ˜1.1 (Sections 3.3 and 3.4) and provide axioms for $\mathcal{L}_{P}$ (Section˜3.5)
Section˜4 concerns Schubert varieties of subspace arrangements. Foundational work of Hassett and Tschinkel [HT99] is reviewed in Sections 4.1 and 4.2. We construct $Y_{V}$ in Section˜4.4, then compute the support of its Chow class in Section˜4.5. In Section˜4.6, we discuss “polymatroid generality”, a condition on $V$ that guarantees $Y_{V}$ consists of finitely many $V$-orbits. Assuming this condition, we identify strata of $Y_{V}$ in Section˜4.7, culminating with a proof of ˜1.3. In Section˜4.8, we prove ˜1.7.
At the last, we pose a large number of questions about Schubert varieties of subspace arrangements in Section˜5. Both algebraic and geometric aspects of Schubert varieties of subspace arrangements seem markedly more complicated than those of Schubert varieties of hyperplane arrangements, yet examples suggest a good theory lies in wait. Interested parties are encouraged to seek their fortune.
@@ 1.6. Acknowledgements
We thank Lucas Gierczak for helping to sort out the relationship between the “matricubes” of [AG24] and the objects of the present work (see ˜3.8), and Matt Larson for helpful comments on our draft. We are all partially supported by the Simons Foundation. The first author is partially supported by NSF grant DMS-2039316, and the last author is partially supported by NSF grant DMS-1926686. We are grateful to the Institute for Advanced Study (IAS). The second and third authors also thank the Beijing International Center for Mathematical Research (BICMR) for their hospitality and support during the preparation of this paper.
@@ 2. Polymatroids
$\{\text{caged polymatroids $(P,\mathbf{n})$}\}$ $\overset{\sim}{\longrightarrow}\{\text{multisymmetric matroids on $\widetilde{E}$}\}$
@@
The matroid lift satisfies $\operatorname{rk}_{P}(S)=\operatorname{rk}_{\widetilde{P}}(\cup_{i\in S}E_{i})$. By applying this fact and ˜2.4, we learn that the closure in $\widetilde{P}$ of $\cup_{i\in S}E_{i}$ is $\cup_{i\in\overline{S}}E_{i}$, where $\overline{S}$ is the closure of $S$ in $P$.
@@
By ˜2.5, $\widetilde{P}$ is the unique multisymmetric matroid on $\widetilde{E}$ that lifts $P$.
@@
$\operatorname{rk}_{\widetilde{P}}(A)=\operatorname{rk}_{\widetilde{P}}(\overline{A})$ $=|\overline{A}\setminus{\overline{A}}^{\mathrm{geo}}|+\operatorname{rk}_{\widetilde{P}}(\overline{A}^{\mathrm{geo}})$ by ˜2.4
@@
Let $\mathbb{K}$ be a field and fix $\mathbf{n}\in\mathbb{Z}_{\geq 0}^{N}$. A linear subspace $V\subset\mathbb{K}^{\mathbf{n}}$ realizes both a polymatroid $P$ on $E=\{1,\ldots,N\}$ and a matroid $M$ on $\widetilde{E}=\{(i,j):1\leq i\leq N,\,1\leq j\leq n_{i}\}$. Different $\prod_{i=1}^{N}\operatorname{GL}(n_{i})$-translates of $V$ may realize different matroids; however, the matroid of a general translate is $\widetilde{P}$, and the polymatroid of any translate is $P$.
@@
With notation as in ˜2.2, if $P$ is realized by $V$, then $P\setminus A$ is realized by $\pi_{E\setminus A}(V)$.
@@
$\operatorname{rk}_{T_{\cup_{i\in S}E_{i}}\widetilde{P}}(\cup_{i\in A}E_{i})$ $=\begin{cases}\operatorname{rk}_{\widetilde{P}}(\cup_{i\in A}E_{i})-1,&\text{if $\operatorname{rk}_{\widetilde{P}}(\cup_{i\in A}E_{i})=\operatorname{rk}_{\widetilde{P}}(\cup_{i\in A\cup S}E_{i})$}\\ \operatorname{rk}_{\widetilde{P}}(\cup_{i\in A}E_{i}),&\text{otherwise}\end{cases}$
$=\begin{cases}\operatorname{rk}_{P}(A)-1,&\text{if $\operatorname{rk}_{P}(A)=\operatorname{rk}_{P}(A\cup S)$}\\ \operatorname{rk}_{P}(A),&\text{otherwise}\end{cases}$
@@
Consequently, $\widetilde{T_{S}P}=T_{\cup_{i\in S}E_{i}}\widetilde{P}$ by ˜2.5. ∎
@@
With notation as in ˜2.2, if $S\subset E$, then $T_{S}P$ is realized by $V\cap\pi^{-1}_{E\setminus S}(H)$, with $H$ a general hyperplane in $\prod_{i\in S}W_{i}$.
@@
We may think of both matroids as being multisymmetric on $\widetilde{E}\setminus j$. Geometric subsets of $\widetilde{E}\setminus j$ are of the form $(\cup_{k\in A}E_{k})\setminus j$, with $A\subset E$. By ˜2.5 and ˜2.7, it suffices to show that
@@
Otherwise, suppose that $\operatorname{rk}_{P}(A)<\operatorname{rk}_{P}(A\setminus i)+\operatorname{rk}_{P}(i)$. In this case, ˜2.4 implies that $\overline{(\cup_{k\in A}E_{k})\setminus j}\supset\cup_{k\in A}E_{k}$ in $\widetilde{P}$. Hence,
@@
We defer discussion of realizing reductions to ˜3.17, where we will see it in action en route to proving ˜1.1(ii).
@@ 3. Combinatorial flats
We introduce the poset of combinatorial flats of a polymatroid $P$. We construct the poset using a chosen lift $\widetilde{P}$ in Section˜3.1, then record the effects of deletion and truncation in Section˜3.2. In Sections 3.3 and 3.4, we establish order-theoretic properties and consider the effects of reduction, culminating in a proof of ˜1.1. Finally, in Section˜3.5, we axiomatize posets of combinatorial flats. This provides a cryptomorphic definition of simple polymatroids that generalizes the cryptomorphism of simple matroids and geometric lattices.
Our results on Schubert varieties of subspace arrangements depend only on Sections 3.1 and 3.2, so geometrically-inclined readers may go directly to Section˜4 after reading these.
Throughout Section˜3, we fix notations:
@@
  • $\operatorname{rk}_{P}(\mathbf{s})$ is the rank of a multiset $\mathbf{s}$, in the sense of Section˜1.2.
@@
A combinatorial flat of $(P,\mathbf{n})$ is a multiset $\mathbf{s}\in\mathbb{Z}_{\geq 0}^{E}$ such that there exists a flat $F$ of $\widetilde{P}$ satisfying $\mathbf{s}=(|F\cap E_{1}|,\ldots,|F\cap E_{N}|)$. (Rephrased in the language of Section˜1.5, $\mathbf{s}$ is a combinatorial flat if it is represented by a flat of $\widetilde{P}$.)
@@
Viewing $\mathcal{L}_{P,\mathbf{n}}$ as a quotient and applying ˜2.4, we learn that rank and closure of multisets have the following properties:
@@
Choose $S\subset\widetilde{E}$ representing $\mathbf{s}$. If $\mathbf{b}$ is a basis of $\mathbf{s}$, then $\mathbf{b}$ is represented by a basis $B$ of $S$ in $\widetilde{P}$ by ˜2.9. The first assertion holds because $\operatorname{rk}_{P,\mathbf{n}}(\mathbf{s})=\operatorname{rk}_{\widetilde{P}}(S)=|B|=|\mathbf{b}|$.
If $\mathbf{s}$ is a combinatorial flat, then $S\setminus S^{\mathrm{geo}}$ consists of coloops of the flat $S$ by ˜2.4. Hence, bases of $S$ are of the form $B^{\prime}\cup(S\setminus S^{\mathrm{geo}})$ with $B^{\prime}$ a basis for $S^{\mathrm{geo}}$. The second assertion follows. ∎
@@
˜3.4 shows that all bases of a multiset $\mathbf{s}$ have the same cardinality, so $\operatorname{rk}_{P}(\mathbf{s})$, as defined in Section˜1.2, is well-defined.
@@
˜3.4 explains the equivalence of the different definitions of “rank” and “combinatorial flat” given in Section˜1.2 and Section˜3.1. The first statement of ˜3.4 says exactly that $\operatorname{rk}_{P,\mathbf{n}}(\mathbf{s})=\operatorname{rk}_{P}(\mathbf{s})$ for any multiset $\mathbf{s}$. When $\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$, this means $\operatorname{rk}_{P}(\mathbf{s})$ is also equal to the rank of $\mathbf{s}$ in the graded poset $\mathcal{L}_{P,\mathbf{n}}$. We will henceforth write only $\operatorname{rk}_{P}(\mathbf{s})$.
For equivalence of the two definitions of “combinatorial flat”, choose $S\subset\widetilde{E}$ representing $\mathbf{s}$. If $s_{i}<n_{i}$, then $\mathbf{s}+\mathbf{e}_{i}$ is represented $S\cup e$ for any $e\in E_{i}\setminus S$. In the sense of Section˜1.2, $\mathbf{s}$ is a combinatorial flat if and only if
@@
for all $e\in E_{i}$ with $s_{i}<n_{i}$. This is precisely the statement that $S$ is a flat of $\widetilde{P}$, which is the notion of “combinatorial flat” defined in Section˜3.1.
@@
Suppose that the polymatroid $P$ is realized by a subspace arrangement $V_{1},\ldots,~V_{N}~\subset~V$. Rephrasing ˜2.10 in the language of arrangements, if $\mathbf{n}$ is a cage for $P$, then $\widetilde{P}$ is realized by any arrangement $\{H_{ij}:1\leq i\leq N,1\leq j\leq n_{i}\}$ in which $H_{i1},\ldots,H_{in_{i}}$ are general hyperplanes containing $V_{i}$. The lattice of flats $\mathcal{L}_{\widetilde{P}}$ has an action by a product of symmetric groups $\mathfrak{S}$, and the quotient $\mathcal{L}_{\widetilde{P}}/\mathfrak{S}$ is isomorphic to $\mathcal{L}_{P}$ (see also Section˜3.1).
@@
From the flag perspective of ˜3.7, the difference between matricubes and combinatorial flats is that matricubes model all collections of flags, while combinatorial flats model only collections of general flags. Combinatorially, this difference manifests in several ways: combinatorial flats obey ˜3.2, are top-heavy, and admit a reasonable notion of basis, but none of these properties hold for matricubes in general. (Failure of top-heaviness and ˜3.2 for matricubes can be seen in the right-hand example of [AG24, Section 3.1]. The difficulties of bases for matricubes are discussed in [AG24, Section 9.1].)
@@
The if direction follows from ˜2.12 and ˜2.13. For the “only if”, suppose that $\pi_{E\setminus i}(\mathbf{s})$ is a combinatorial flat of $(P\setminus i,\pi_{E\setminus i}(\mathbf{n}))$. We will show that $\mathbf{s}$ must be a flat by proving $\mathbf{s}=\overline{\mathbf{s}}$. By ˜3.2 and the hypothesis that $\overline{s}_{i}<n_{i}$, it is always the case that $s_{i}=\overline{s}_{i}$. Hence, it suffices to show that $s_{j}=\overline{s}_{j}$ for $j\neq i$.
By ˜2.12 and ˜2.13, there is a combinatorial flat $\mathbf{s}^{\prime}$ of $(P,\mathbf{n})$ with $\pi_{E\setminus i}(\mathbf{s}^{\prime})=\pi_{E\setminus i}(\mathbf{s})$. If $s^{\prime}_{i}=n_{i}$, then $\mathbf{s}<\mathbf{s}^{\prime}$, so $\mathbf{s}\leq\overline{\mathbf{s}}<\mathbf{s}^{\prime}$. Entrywise, this implies $s_{j}\leq\overline{s}_{j}\leq s_{j}^{\prime}=s_{j}$, which gives the desired equality.
Otherwise, $s_{i}^{\prime}<n_{i}$. By Lemmas 3.2 and 3.3, $\mathbf{s}^{\prime\prime}:=\mathbf{s}^{\prime}-s_{i}^{\prime}\mathbf{e}_{i}=\mathbf{s}-s_{i}\mathbf{e}_{i}$ is a combinatorial flat of rank at least $\operatorname{rk}(\mathbf{s})-s_{i}$. By ˜3.2, $\overline{\mathbf{s}}-s_{i}\mathbf{e}_{i}$ is also a combinatorial flat, and has rank $\operatorname{rk}(\overline{\mathbf{s}})-s_{i}$. Since $\mathbf{s}^{\prime\prime}\leq\overline{\mathbf{s}}-s_{i}\mathbf{e}_{i}$, these two combinatorial flats must be equal. In particular, for $j\neq i$, $s_{j}^{\prime\prime}=\overline{s}_{j}$. Since $s_{j}^{\prime\prime}=s_{j}$ by construction, this means $s_{j}=\overline{s}_{j}$, as desired. ∎